### v1

@@ Appendix D Difference-in-Differences
Goodman-Bacon (2021) provides the following representation of the two-way fixed effects estimand under the assumption that group-level average treatment effects are constant over time:
@@
$\beta_{\text{TWFE}}$ $=$ $\frac{\mathbb{E}[a_{B}(G)\cdot\mathbb{P}(D=1\mid G)\cdot\mathbb{E}[Y(1)-Y(0)\mid G,D=1]]}{\mathbb{E}[a_{B}(G)\cdot\mathbb{P}(D=1\mid G)]}$
$=$ $\frac{\sum_{k:\;\operatorname{var}(D\mid G=k)>0}\mathbb{P}(G=k)\cdot a_{B}(k)\cdot\mathbb{P}(D=1\mid G=k)\cdot\mathbb{E}[Y(1)-Y(0)\mid G=k,D=1]}{\sum_{k:\;\operatorname{var}(D\mid G=k)>0}\mathbb{P}(G=k)\cdot a_{B}(k)\cdot\mathbb{P}(D=1\mid G=k)},$
it becomes clear that, for each group $k$ other than the always treated and the never treated,
$a_{B}(k)\cdot\mathbb{P}(D=1\mid G=k)$ $=$ $\sum_{j=1}^{k-1}\mathbb{P}(G=j)\cdot\mathbb{P}(D=1\mid G=k)\cdot\big{[}\mathbb{P}(D=1\mid G=j)-\mathbb{P}(D=1\mid G=k)\big{]}$
@@
$a_{B}(k)$ $=$ $\sum_{j=1}^{k-1}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=j)-\mathbb{P}(D=1\mid G=k)\big{]}$ (D.1)
$+$ $\sum_{j=k+1}^{K}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=k)-\mathbb{P}(D=1\mid G=j)\big{]}\cdot\frac{1-\mathbb{P}(D=1\mid G=k)}{\mathbb{P}(D=1\mid G=k)}.$
D.1 Equivalence of Weight Functions
We now show that the weights obtained in equation (D.1) are equivalent to those in Proposition 2.5. First, we rewrite the weights in (D.1) as follows:
$a_{B}(k)$ $=\sum_{j=1}^{k-1}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=j)-\mathbb{P}(D=1\mid G=k)\big{]}$
@@ D.1 Equivalence of Weight Functions
$=(\mathbb{E}[D]-\mathbb{E}[D\mid G=k]\mathbb{P}(G=k))-\mathbb{E}[D\mid G=k]\mathbb{P}(G\neq k)+\mathbb{P}(G>k)\left(1-\frac{\mathbb{E}[D\mid G>k]}{\mathbb{E}[D\mid G=k]}\right)$
@@
$\mathbb{E}[1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D]\mid G=k]$ $=1-\mathbb{E}[D\mid G=k]-\mathbb{E}[D\mid P\geq k]+\mathbb{E}[D].$ (D.2)
@@
$\mathbb{E}[1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D]\mid G=k]$ $=\mathbb{P}(D=0\mid G=k)\cdot(\mathbb{P}(D=0\mid P\geq k)+\mathbb{P}(D=1\mid P<g)).$
Let us look at the difference between the weights in (D.1) and (D.2). Fix $k\in\{2,\ldots,T\}$ and write
$\left(1-\mathbb{E}[D\mid G=k]-\mathbb{E}[D\mid P\geq k]+\mathbb{E}[D]\right)-\left(\mathbb{E}[D]-\mathbb{E}[D\mid G=k]+\mathbb{P}(G>k)\left(1-\frac{\mathbb{E}[D\mid G>k]}{\mathbb{E}[D\mid G=k]}\right)\right)$
@@
$=\frac{1}{\mathbb{E}[\mathbbm{1}(k\leq P)]}\left(F_{G}(k)\mathbb{E}[\mathbbm{1}(k\leq P)]+\mathbb{E}[F_{G}(P)\mathbbm{1}(P\geq k)]-F_{G}(k)\mathbb{E}[\mathbbm{1}(P\geq k)]-\mathbb{E}[F_{G}(P)\mathbbm{1}(P\geq k)]\right)$
@@
Therefore, the weights in Proposition (2.5), (D.1), and (D.2) are all equal to one another.

### v3

@@ Appendix H Additional Derivations for Difference-in-Differences
Goodman-Bacon (2021) provides the following representation of the two-way fixed effects estimand under the assumption that group-time average treatment effects are constant over time:
@@
$\beta_{\text{TWFE}}$ $=$ $\frac{\mathbb{E}[a_{\text{TWFE,H}}(G)\cdot\mathbb{P}(D=1\mid G)\cdot\tau_{0}(G)]}{\mathbb{E}[a_{\text{TWFE,H}}(G)\cdot\mathbb{P}(D=1\mid G)]}$
$=$ $\frac{\sum_{k:\;\operatorname{var}(D\mid G=k)>0}\mathbb{P}(G=k)\cdot a_{\text{TWFE,H}}(k)\cdot\mathbb{P}(D=1\mid G=k)\cdot\tau_{0}(k)}{\sum_{k:\;\operatorname{var}(D\mid G=k)>0}\mathbb{P}(G=k)\cdot a_{\text{TWFE,H}}(k)\cdot\mathbb{P}(D=1\mid G=k)},$
where $\tau_{0}(k)=\mathbb{E}[Y(1)-Y(0)\mid G=k,D=1]$, it becomes clear that, for each group $k$ other than the always treated and the never treated,
$a_{\text{TWFE,H}}(k)\cdot\mathbb{P}(D=1\mid G=k)$
$=\sum_{j=1}^{k-1}\mathbb{P}(G=j)\cdot\mathbb{P}(D=1\mid G=k)\cdot\big{[}\mathbb{P}(D=1\mid G=j)-\mathbb{P}(D=1\mid G=k)\big{]}$
@@
$a_{\text{TWFE,H}}(k)$
$=\sum_{j=1}^{k-1}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=j)-\mathbb{P}(D=1\mid G=k)\big{]}$
$+\sum_{j=k+1}^{K}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=k)-\mathbb{P}(D=1\mid G=j)\big{]}\cdot\frac{1-\mathbb{P}(D=1\mid G=k)}{\mathbb{P}(D=1\mid G=k)}.$ (A.1)
Equivalence of Weight Functions
We now show that the weights obtained in equation (A.1) are equivalent to those in Proposition 5.4. First, we rewrite the weights in (A.1) as follows:
$a_{\text{TWFE,H}}(k)$
$=\sum_{j=1}^{k-1}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=j)-\mathbb{P}(D=1\mid G=k)\big{]}$
@@
$=(\mathbb{E}[D]-\mathbb{E}[D\mid G=k]\mathbb{P}(G=k))-\mathbb{E}[D\mid G=k]\mathbb{P}(G\neq k)$
$\quad+\mathbb{P}(G>k)\left(1-\frac{\mathbb{E}[D\mid G>k]}{\mathbb{E}[D\mid G=k]}\right)$
@@
$\mathbb{E}[1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D]\mid G=k]$ $=1-\mathbb{E}[D\mid G=k]-\mathbb{E}[D\mid P\geq k]+\mathbb{E}[D],$ (A.2)
@@
$\mathbb{E}[1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D]\mid G=k]$
$=\mathbb{P}(D=0\mid G=k)\cdot(\mathbb{P}(D=0\mid P\geq k)+\mathbb{P}(D=1\mid P<g)).$
Let us look at the difference between the weights in (A.1) and (A.2). Fix $k\in\{2,\ldots,T\}$ and write:
$\left(1-\mathbb{E}[D\mid G=k]-\mathbb{E}[D\mid P\geq k]+\mathbb{E}[D]\right)$
$\quad-\left(\mathbb{E}[D]-\mathbb{E}[D\mid G=k]+\mathbb{P}(G>k)\left(1-\frac{\mathbb{E}[D\mid G>k]}{\mathbb{E}[D\mid G=k]}\right)\right)$
@@
$=\frac{1}{\mathbb{E}[\mathbbm{1}(k\leq P)]}\left(F_{G}(k)\mathbb{E}[\mathbbm{1}(k\leq P)]+\mathbb{E}[F_{G}(P)\mathbbm{1}(P\geq k)]-F_{G}(k)\mathbb{E}[\mathbbm{1}(P\geq k)]\right.$
$\left.\quad-\mathbb{E}[F_{G}(P)\mathbbm{1}(P\geq k)]\right)$
@@
Therefore, the weights in Proposition 5.4 and equations (A.1) and (A.2) are all equal to one another.